OpenStudy (anonymous):

describe how two different ways to use elimination to solve the system {3x-4y=-1 5x+4y=-5

OpenStudy (anonymous):

*how to use

OpenStudy (anonymous):

first try adding the two equations together: 3x - 4y = -1 + (5x + 4y) = -5

OpenStudy (anonymous):

okay